OpenStudy (anonymous):

Riley noticed an aquarium in his father’s office. The aquarium had the dimensions 16 in. by 8.5 in. by 10.5 in. The formula for volume is: V = l × w × h. (a) Riley noticed that the aquarium was three-fourths full of water. How many more cubic inches of water would be required to fill the tank?

OpenStudy (anonymous):

@jabez177

OpenStudy (anonymous):

i have no idea what it is

jabez177 (jabez177):

Hmmm... I will try my best to solve this equation.

OpenStudy (anonymous):

the volume is \[1428\]

OpenStudy (anonymous):

thats all I know

jabez177 (jabez177):

Try multiplying that by .75

OpenStudy (anonymous):

\[1428*0.75=1071\]

jabez177 (jabez177):

Now subtract 1071 from 1428 "1428-1071"

OpenStudy (anonymous):

\[1428-1071=357\]

jabez177 (jabez177):

Maybe that is the answer because it wants to know how much more water is needed to fill up the tank right? Math is hard. :P

OpenStudy (anonymous):

yes i will put that in. and this is part (a) so can i ask you part (b) and (c) as well?

jabez177 (jabez177):

Okay. I don't want to get you wrong though. :(

OpenStudy (anonymous):

(b) Another aquarium in the building has dimensions that are each triple the size of the aquarium in Riley’s father’s office. Riley thought that the volume would also triple. Is Riley correct? How many times greater is the volume of the larger aquarium than the smaller? Show your work and explain your reasoning.

OpenStudy (anonymous):

would i do 16*3, 8.5*3, and 10.5*3?

jabez177 (jabez177):

She is correct, but hold on, i'll be right back

OpenStudy (anonymous):

k

OpenStudy (anonymous):

are you back yet

jabez177 (jabez177):

Now I am, I'm eating. :P

OpenStudy (anonymous):

ok but you were wrong i did the math this is the right answer but thanks anyway. She is incorrect because if you triple all of the dimensions the volume will not be tripled, it would equal a different number.. 16*3=48, 8.5*3=25.5,10.5*3=31.5. 48*25.5*31.5=38,556. 1428*3=4284

jabez177 (jabez177):

Wow! You are smarter than me. :P I just haven't had a lot of practice with these in a Loooooong time. :P Nice job! :)

OpenStudy (anonymous):

ok thx thats all i have i thought i had a letter (c) to this question but nope thanks for your help